the exponential model describes the population, A, of India, in millions, t years after 1992. A=1072.2e^.012t In which year will India's population be 1653 million?
1
Expert's answer
2020-04-27T18:14:52-0400

Given,


A = "1653 \space"millions

We know, one million = "10^6"


A = "1653 \times 10^6"


t = number of years

"A = 1072.2e^{.012t}"



"1653 \times 10^6 = 1072.2 \times e^{0,012\times t}"



"e^{0.012 \times t} = \frac {1653 \times 10^6}{1072.2}"

"e^{0.012 \times t} = 1,541,689.98"

Apply log on both sides


"ln (e^{0.012 \times t}) = ln (1,541,689.98)"

"0.012 \times t = 14.24"

"t = \frac {14.24}{0.012} =1186.66 \approx 1187"

After "1187" years, the population will becomes "1653" millions


Hence, year = "1992 + 1187 =3179"